OpenStudy (anonymous):

a survey of 500 television watchers produced the information: 285 watch football games 195 watch hockey games 115 watch basketball games 45 watch football and basketball 70 watch football and hockey 50 watch hockey and basketball 50 do not watch any of the three games A. how many people in the survey watch all the three games? B. How many people watch exactly one of the three games?

OpenStudy (anonymous):

In venn diagram

OpenStudy (anonymous):

The principle of inclusion-exclusion would take care of this...but im a bit fuzzy on the finer details....one sec, going through old notes lol

myininaya (myininaya):

i think A is 310

myininaya (myininaya):

we have 500 people the sum of number of elements in each set total 810

myininaya (myininaya):

so we have some overlapping

myininaya (myininaya):

does anyone think anything of my answer to A

myininaya (myininaya):

maybe i should have done 810-450

OpenStudy (anonymous):

im checking it right now, actually, let me upload the rules for the inclusion exclusion principle with some examples, you could probably understand it a little better than i can lol

myininaya (myininaya):

i will look it up

OpenStudy (anonymous):

im getting 20 for the answer lol

OpenStudy (anonymous):

for the number of people that watch all 3

OpenStudy (anonymous):

THE NUMBER OF PEOPLE WHO WATCH ALL 3 IS 10.

OpenStudy (anonymous):

its: 500-285-195-115+70+45 +50 - 50

OpenStudy (anonymous):

im 100% sure its 20 now

OpenStudy (anonymous):

Heres how Inclusion Exclusion works: Let N(a) be the number of people that watch football, N(b) for hockey, N(c) for basketball. N(ab) for football and hockey, N(ac) for football and basketball, N(bc) for hockey and basketball, N(abc) for all three (what we want), and N(a'b'c') for those that watch none of those three.

OpenStudy (anonymous):

The formula is: N(a'b'c') = (Total Number of people) -N(a)-N(b)-N(c)+N(ab)+N(ac)+N(bc)-N(abc)

myininaya (myininaya):

football watchers=F Hockey watchers=H Basketball watchers=B |F n H n B|=|F|+|H|+|B|-|F U H U B| -|F n B|- |F n H|-|H n B| =285+195+115-|F U H U B| -70-45-50 =430-|F U H U B| what is |F U H U B|=?

OpenStudy (anonymous):

so we end up with: 50 = 500-285-195-115+70+45+50- N(abc)

OpenStudy (anonymous):

solving for N(abc) gives 20

OpenStudy (anonymous):

285+195+115-450=145 WHO ARE THE COMMON WATCHERS OF DIFFERENT GAMES. 50-X+X+70-X+45-X=145 X=10

OpenStudy (anonymous):

This is the rules of the principle, uploading some examples in a sec.

OpenStudy (anonymous):

er, an example. im still 100% sure the answer is 20, its a straightfoward application of this.

myininaya (myininaya):

here i think found an example havent read through it but i think it will help http://www.math.uic.edu/~fye/Math118Fall2010/Practice-Final-with-Answers.pdf

myininaya (myininaya):

no i dont like that example that i posted

OpenStudy (anonymous):

i can answer c of that example, but not the other 2, at least in a straightforward manner anyways

myininaya (myininaya):

ok i think i agree with joe he showed proof for his formula and i believe i get it

OpenStudy (anonymous):

Thank you :) after that spherical triangle mess i was thinking i wasnt going to get through to anyone anymore ;; i still dont know how to answer b of this question yet though lol >.<

myininaya (myininaya):

so what is the not exactly one

myininaya (myininaya):

at least one right?

myininaya (myininaya):

i think im thinking of probability

myininaya (myininaya):

remember this P(C)=1-P(not C)

OpenStudy (anonymous):

right right

myininaya (myininaya):

ok but i know we arent talking about probability but lets think C1 and C2 are events of C P(C)=1 P(C1)=P(C)-P(not C1) so does this hold C1=C-notC1

OpenStudy (anonymous):

im gonna stare at that example you posted earlier, i think that would help

myininaya (myininaya):

for exactly two in there example they got 2+11+1 let me think where did they get the numbers

myininaya (myininaya):

i think i see it

myininaya (myininaya):

(285-20)+(195-20)+(115-20)

OpenStudy (anonymous):

20 comes from our answer from a right?

myininaya (myininaya):

nope thats not right lol

myininaya (myininaya):

well they have (7-5)+(16-5)+(6-5) right? let me look back

myininaya (myininaya):

yes joe 20 came from our answer above

myininaya (myininaya):

the info for all number who enjoyed all 3 was given to us in the example and 7,16,6 were the people who enjoyed both of something

myininaya (myininaya):

the number that enjoyed all was 5

OpenStudy (anonymous):

so 7-5, 16-5, and 6-5 gave the number of people that enjoyed only 2 activities

myininaya (myininaya):

right!

myininaya (myininaya):

i was think if we took it away from the groups who enjoyed 1 thing we would have it

OpenStudy (anonymous):

so 2, 11, and 1, or 14 people enjoyed only 2, 5 people enjoyed 3, so how can we use this to figure out how many people only enjoyed one?

OpenStudy (anonymous):

19 people didnt enjoy anything (used the formula above to get that)

OpenStudy (anonymous):

ah i think i got it

OpenStudy (anonymous):

83 people surveyed. Lets take out the number of people that liked nothing, exactly 2 things, or 3 things. that should leave the number of people that only like exactly one thing. 83 - 19 - 14 - 5 = 45 which is the sum of the numbers in that venn diagram of people that only liked one sport, 15+17+13

OpenStudy (anonymous):

So now to apply it to this problem....

OpenStudy (anonymous):

it would be the 500 people total, minus the people that watch nothing, the people that watch everything, and the people that watch only 2 things

OpenStudy (anonymous):

we gotta calculate the number of people that only watch 2 things, but we figured that out already earlier, its: 70-20+45-20+50-20 = 105

OpenStudy (anonymous):

so the number of people that only watch one thing is: 500-50-20-105 = 325

OpenStudy (anonymous):

So.....after all this deliberation, im sticking with a) 20 , b) 325

OpenStudy (anonymous):

how can i get 20??

myininaya (myininaya):

hey you might like this site prince http://interactives.mped.org/view_interactive.aspx?id=28&title=

OpenStudy (anonymous):

>.> lol let x be the number of people that watch all 3, then you need to solve this equationL 50 = 500 - 285 - 195-115+70+45+50-x

OpenStudy (anonymous):

ooo that is cool lol

myininaya (myininaya):

gj joe im still not sure about b

myininaya (myininaya):

but joe is really smart so he is probably right

myininaya (myininaya):

he has probability of 91/100 of being right

myininaya (myininaya):

oh what i just got 325

myininaya (myininaya):

i drew a venn diagram and played it with and i did 190+40+95

OpenStudy (anonymous):

nice nice, does this confirm the answers then?

myininaya (myininaya):

i have 190 only watch football i have 40 only watch basketball i have 95 only watch hockey i have 20 only watch all 3 i have 50 only watch both Hockey and football i have 30 only watch both basketball and and hockey i have 25 only watch both fooball and basketball

OpenStudy (anonymous):

now we could answer any question lol, nice job :)

myininaya (myininaya):

yep yep

myininaya (myininaya):

awesome job joe

OpenStudy (anonymous):

another application of inclusion exclusion: how many numbers from 1 to 100 arent divisible by 2, 3, or 5?

myininaya (myininaya):

this was one of your projects?

OpenStudy (anonymous):

i went to a seminar on it and got that packet i posted (had some other examples and problems). i dont know it there is a class that uses this, maybe a combinatorics class? since it involves counting things in an efficient manner?

OpenStudy (anonymous):

This deserves breakfast, im gonna go get breakfast <.< i suggest you all do the same lol.